OpenStudy (anonymous):

Tricky Algebra 2 problem. :/ It's solving for extraneous solutions. Take a look please?

OpenStudy (anonymous):

\[9\left| 9-8x \right| = 2x + 3\]

OpenStudy (anonymous):

So far I have \[9 - 8x = \frac{ 2 }{ 9}x + \frac{ 1 }{ 3}\]

OpenStudy (anonymous):

so your equation is good.. find x. but there is another equation that satisfies theequation in the question

OpenStudy (anonymous):

would i subtract 9 from both sides? and i will do the other equation too :)

OpenStudy (anonymous):

so we take all the terms without x to one side and the terms with x to the other. you can subtract 9 from both sides and then subtract also 2x/9 from both sides.

OpenStudy (anonymous):

So after subtracting 9 from both sides I get \[8x = \frac{ 2 }{ 9}x -\frac{ 26 }{ 3}\]

OpenStudy (anonymous):

i wrote: |x| = a gives two options : x=a and x=-a there should be an "or" instead of the "and" between the solutions :)

OpenStudy (anonymous):

then i would subtract 2/9x from both sides. right?

OpenStudy (anonymous):

what happened to the - before the 8x ?

OpenStudy (nurali):

9|9−8x|=2x+3 |9−8x|=2x+3/9 9−8x=2x+3/9 or 9−8x=-(2x+3/9) 9(9−8x)=2x+3 or 9(9−8x)=-(2x+3) 81-72x=2x+3 or 81-72x=-2x-3 81-3=2x+72x or 81+3=-2x+72x 78=74x or 84=70x x=78/74 or x=84/70 x=39/37 or x=6/5

OpenStudy (anonymous):

oops almost left that out!

OpenStudy (anonymous):

@Nurali why did you solve it ? you see that she is trying!

OpenStudy (anonymous):

thank you both for helping me :)

OpenStudy (nurali):

My Pleasure.